OpenStudy (anonymous):

solve for x 64=32^(x-4)

OpenStudy (anonymous):

ln64=(x-4)ln32 then (ln64/ln32)=x-4 then (6ln2/5ln2)=x-4 then (6/5)+4 =x therefore x=26/5